OpenStudy (anonymous):

Kelly makes $10 more per hour than her regular pay when she is working overtime. Yesterday, she worked 8 regular hours and 4 overtime hours and earned $160. What is Kelly's regular pay per hour? $20 per hour $10 per hour $12 per hour $16 per hour

OpenStudy (whpalmer4):

regular pay = x overtime = x+10 8 hours regular + 4 overtime = 160 8x + 4(x+10) = 160 Solve for x

OpenStudy (triciaal):

LET X = REG THEN OVERTIME =10 + X 8X +4(10 + X )= 160 8X +40 + 4X =160 12X =160 - 40 X =10